hannieee:

The ratio of the ages of Mandy and Sandy is 2:5. After 8 years, their ages will be in the ratio 1:2. What is the difference between their present ages? Caitlin says the difference in their ages is 24 years. In two or more complete sentences, explain whether or not Caitlin is correct. In your explanation, include the equations and calculations necessary to support or contradict Caitlin's answer.

hannieee:

\[\frac{ m+8 }{ s+8}=\frac{ 1 }{ 2 }\] This is as far as i have been able to get

hannieee:

apparently the next step is 2m + 16 = s + 8 2m + 8 = s 2m + 8 = But i don't understand where the 16 is coming from

Extrinix:

Cross multiplication, \(2(m+8)\) = \(1(s+8)\)

jhonyy9:

@hannieee wrote:
apparently the next step is 2m + 16 = s + 8 2m + 8 = s 2m + 8 = But i don't understand where the 16 is coming from
here are what you wrote above @extrinix

hannieee:

i dont think that -8 is the correct answer though

hannieee:

alright. and then what?

hannieee:

i got 2m+8=s

hannieee:

m=s/2-4

hannieee:

?

hannieee:

ohhh! Okay. And where do i go from here?

hannieee:

okay. How should i do that?

Extrinix:

skrrt, here's an easy way around this. 971847

hannieee:

????

Extrinix:

There's an explanation that's morein depth on that link.